A line that passes through (8,4) and is perpendicular to 2x-y=-7
Answer by lwsshak3(11628)

A line that passes through (8,4) and is perpendicular to 2x-y=-7
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
2x-y=-7
y=2x+7
slope=2
For line passing thru (8,4) and perpendicular to first equation
slope=-1/2 (negative reciprocal)
equation: y=-x/2+b
using given point (8,4) to solve for b
4=-8/2+b
b=8
Equation of perpendicular line:
y=-x/2+8